OpenStudy (tatumlee):

Evaluate the function for x=2:f(x)=7x-5 can someone show me how to do this please?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=7x-5 }\) This function gives you the output \(f(x)\) for every value of \(x\) that you plug into the function. Here, I will give you an *example*. If I wanted to evaluate the function for *x=3*, then this is what I would do: \(\large\color{black}{ \displaystyle f(3)=7\cdot3-5 }\) \(\large\color{black}{ \displaystyle f(3)=21-5 }\) \(\large\color{black}{ \displaystyle f(3)=16 }\)

OpenStudy (solomonzelman):

You need to evaluate the function for *x=2* (not x=3). That means thaty you need to plug *2* instead of *x* into the function f(x). (Just as I did with 3, above, -- but this time with a "2")

OpenStudy (tatumlee):

Would the answer be 9? I did 7 x 2-5 which is 9

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(2)=7\cdot2-5=14-5=9 }\)\ You are exactly right! Good job!

OpenStudy (tatumlee):

Thank you so much!

OpenStudy (solomonzelman):

yw